If a book is 80 square inches How many 1x1x1 centimeter trianles wwould it take to cover the book? Each side of the triangle is 1 centimeter and 2.54 centimeters equal one inch.
Answer by checkley71(8403)

EACH TRINGLE HAS AN AREA OF
1/2B*H OR .5*1*H AND THE HEIGHT IS FOUND BY THE PATHAGORIAN THE0REM A^2+B^2=C^2.
WE HAVE C=1 & B=.5 NOW WE SOLVE FOR A
A^2+.5^2=1^2
A^2+.25=1
A^2=1-.25
A^2=.75 NOW WE CALCULATE THE SQUARE ROOT OF BOTH SIDES & WE GET
A=.866 FOR THE HEIGHT. NOW WE CAN CALCULATE THE AREA OF EACH TRIANGLE IN CMs.
AREA=.5*1*.866
AREA=.433 SQUARE CMs.
NOW WE CAN CONVERT THE SIZE OF THW BOOK IN INCHES TO CM OR CONVERT THE TRIANGLE IN CMs TO INCHES. LETS CONVERT THE 80 SQ. INCHES TO CMs.
80*2.54=203.2 SQ CMs.
NOW WE DIVIDE THE BOOK AREA BY THE SIZE OF THE TRIANGLE THUS
203.2/.433=469.284 TRIANGLES WOULD COVER THE BOOK.